# Universal Predictors for Mixing Time more than Liouvillian Gap

### v1

@@ Abstract
A key concept in the study of open system dynamics is the mixing time, defined as the minimum time it takes for an arbitrary initial state to evolve to the system’s steady state [54, 79]. In dissipative state preparation, mixing time is a vital metric because it quantifies the practical time required to prepare a desired state, thus determining whether a protocol is experimentally feasible [18, 74]. Also, an estimation of the mixing time typically plays a crucial role in many of the quantum algorithmic applications to yield efficient quantum algorithms [28, 1]. Also, mixing time is related to entropy [31, 46, 13], entanglement dynamics, and the structure of the Lindblad spectrum and the dissipative phase transition, where at the transition point, the mixing time diverges [46, 79, 45, 42, 57, 33, 32].
Lots of progress has been made in analyzing the mixing time for the open system dynamics. Previous studies have shown that for Lindbladians satisfying quantum detailed balance, the system’s dynamics can be mapped to a local, frustration-free parent Hamiltonian in a doubled Hilbert space. The spectral gap of this Hamiltonian then provides a useful lower bound for the Lindbladian’s spectral gap, offering a powerful tool for estimating mixing time [15, 68, 35]. A different approach uses the modified logarithmic Sobolev inequality [73, 9], which analyzes the exponential decay of relative entropy. This method can provide a tighter, polylog($N$) mixing time bound, offering a significant improvement over bounds based solely on the spectral gap [36, 11, 4, 3, 23]. Besides, other research has estimated mixing time by analyzing the Hamiltonian and dissipative parts separately, although this is applicable in limited cases [20]. Despite the recent progress, obtaining the mixing time of the Lindbladian dynamics, and analyzing its dependent of system size is typically difficult and can sometimes be the bottleneck of the analysis.
@@
That is, $\tau_{\mathrm{mix}}(\eta)$ is the minimal time such that, for arbitrary initial state $\rho_{0}$, the evolved state $e^{\mathcal{L}t}[\rho_{0}]$ is within trace distance $\eta$ of the steady state $\sigma$. Since density matrices are Hermitian, this distance can also be written as
$D(\rho(t),\sigma)=\frac{1}{2}\,\mathrm{Tr}\,|\rho(t)-\sigma|.$ (4)
@@
Here $L$ and $R$ denote the left and right copies, respectively. Under this mapping, the Lindblad master equation becomes a Schrödinger-like equation in doubled space [76],
@@
$\begin{split}H_{s}=&H_{L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes H_{R}^{T},\\ H_{d}=&\gamma\sum_{\alpha}\Big[-2\,K_{\alpha,L}\otimes K_{\alpha,R}^{*}\\ &+(K_{\alpha}^{\dagger}K_{\alpha})_{L}\otimes\mathcal{I}_{R}+\mathcal{I}_{L}\otimes(K_{\alpha}^{\dagger}K_{\alpha})_{R}^{*}\Big].\end{split}$ (7)
@@
where $\epsilon_{j}=\alpha_{j}-i\beta_{j}$ are the eigenvalues of $H^{D}$ (the Lindblad spectrum). The real parts $\alpha_{j}$ generate oscillations, while $\beta_{j}\geq 0$ set decay rates. [^1] Complete positivity implies $\beta_{j}\geq 0$. Trace preservation fixes the steady mode: $c_{0}=1$, $\mathrm{Tr}\,\sigma_{0}=1$, and $\mathrm{Tr}\,\sigma_{j}=0$ for $j\neq 0$. Hermiticity preservation implies $\sigma_{j}=\sigma_{j}^{\dagger}$.
Consequently, the deviation from the steady state $\sigma_{0}$ is
$|\psi_{\rho}^{D}(t)\rangle-|\psi_{\sigma_{0}}\rangle=\sum_{j\neq 0}c_{j}\,e^{-i\alpha_{j}t-\beta_{j}t}\,|\psi_{\sigma_{j}}\rangle,$ (10)
@@
$|\psi_{\rho_{0}}\rangle=|\psi_{\sigma_{0}}\rangle+c_{1}|{\psi_{\sigma_{1}}}\rangle,$ (11)
where $\sigma_{0}$ is the steady state and $\sigma_{1}$ is the slowest decaying mode of the Lindbladian with eigenvalue $\Delta+i\alpha_{1}$ ($\Delta>0$). One might wonder whether the coefficient $c_{1}$ could grow with system size. This is ruled out by a purity bound: under vectorization, $\|\lvert\psi_{\rho_{0}}\rangle\|^{2}=\mathrm{Tr}(\rho_{0}^{2})\leq 1$, so the overlap (and hence $c_{1}$) is at most $O(1)$. We may therefore treat $c_{1}$ as an $O(1)$ constant.
At time $t$,
$|\psi_{\rho_{0}}(t)\rangle-|\psi_{\sigma_{0}}\rangle=c_{1}e^{-(\Delta-i\alpha_{1})t}\,|\psi_{\sigma_{1}}\rangle.$ (12)
Consequently, the trace distance to the steady state is
$D(\rho(t),\sigma_{0})=\frac{|c_{1}|}{2}\operatorname{Tr}|e^{-i\alpha_{1}t-\Delta t}\sigma_{1}|=\frac{|c_{1}|}{2}e^{-\Delta t}\operatorname{Tr}|\sigma_{1}|.$ (13)
@@
$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log(\operatorname{Tr}|\sigma_{1}|)-\log{(\frac{2\eta}{|c_{1}|})}\right].$ (14)
This reformulation of the mixing time indicates that the mixing time of an open system is set by the inverse Liouvillian gap, multiplied by a logarithmic factor controlled by the trace norm of the lowest excited state of the doubled-space Hamiltonian. One might worry that higher excited states, despite their larger decay rates, could have sufficiently smaller trace norms to yield a larger overall mixing-time bound. We explain in Appendix C why it is nevertheless sufficient to focus on the lowest excited state  [^2] We focus on the lowest excited state for two reasons. First, the fast mixing condition imposes no explicit constraints on trace norms across the spectrum, so the lowest excited state provides a natural and minimal diagnostic. Second, in both the weak- and strong-dissipation regimes, the rapid mixing requirement on the trace norm takes the same functional form for all excited states, and is therefore not specific to the lowest excited state. See Appendix C for details..
@@
$\tau_{\textit{mix}}(\eta)\leq\frac{1}{2}\Delta^{-1}\left[\log(N)-2\log{(\frac{2\eta}{|c_{1}|})}\right].$ (17)
@@
$\tau_{\text{mix}}(\eta)\leq\tfrac{1}{2}\Delta^{-1}\left[L\log d_{0}-2\log{(\frac{2\eta}{|c_{1}|})}\right].$ (18)
This bound of mixing time is general but not always saturated. For example, non-Hermitian skin effects can localize eigenstates near boundaries [29], reducing effective support and giving $\tau_{\text{mix}}\sim\Delta^{-1}\text{poly}(\log L)$ for total system size $L$. Mixing time is therefore governed jointly by the Liouvillian gap and the spatial structure of eigenstates.
@@
We next specialize these general criteria to the weak- and strong-dissipation regimes, which yield concrete model-dependent sufficient conditions. The resulting rapid mixing conditions are summarized in Table 2.
@@
Strong-dissipation regime— In this regime, the dissipative part of the Liouvillian dominates, and the Hamiltonian (Schrödinger) contribution can be treated perturbatively. We use degenerate perturbation theory to estimate the Liouvillian gap and the mixing time, and obtain the following fast-mixing corollary:
@@
$\langle\psi_{0k}|V^{D}|\psi_{1\alpha}\rangle\sim J,\qquad\langle\psi_{0k}|V^{D}|\psi_{0l}\rangle\sim J^{\prime},$
where $\{|\psi_{0k}\rangle\}$ spans the degenerate zeroth-order manifold (chosen to diagonalize $V^{D}$ within that subspace) and $\{|\psi_{1\alpha}\rangle\}$ denotes zeroth-order excited modes.
@@
The diagonal shift $V^{D}_{0k,0k}$ contributes only to the real part, while the Liouvillian gap is set by the dissipative contribution from the second-order term,
$\Delta\equiv\mathrm{Im}\,E_{0k}\sim\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{1q}-E^{(0)}_{0k}}\sim c\,\frac{J^{2}}{\gamma_{\mathrm{loc}}},$ (23)
with an $\mathcal{O}(1)$ constant $c$ (for our boundary coupling structure, c=8c=8). [^4] The factor $8$ counts: (i) action on the left or right copy, (ii) action at the left or right boundary, and (iii) raising or lowering the boundary state, giving $2^{3}=8$ possibilities. Crucially, $\Delta$ does not shrink with $L$: only Hamiltonian terms involving boundary sites connect the steady manifold to decaying modes, so the number of contributing excited states is set by the boundary size, i.e., $\mathcal{O}(1)$ in 1D.
@@
$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr).$ (24)
@@
and this bound is saturated here because the $V^{D}$-diagonal basis typically forms superpositions across the full manifold. Therefore,
$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}(\log d_{0})\,L,$ (25)
which is polynomial in $L$ and hence implies fast mixing. In the Appendix we present the analogous analysis for strong bulk dissipation and show that this perturbative mechanism does not apply there.
Physically, boundary-driven fast mixing is a quantum-Zeno-type stabilization: strong edge dissipation sets an $\mathcal{O}(1)$ scale, while boundary Hamiltonian couplings generate a size-independent gap $\Delta\propto J^{2}/\gamma_{\mathrm{loc}}$. Thus, boundary dissipation can guarantee scalable mixing even for complex bulk dynamics, whereas bulk dissipation typically requires different control assumptions (see Appendix).
Finally, in the strong-dissipation regime, rapid mixing also requires control of the trace norm $\textrm{Tr}|\sigma_{1}|$, as stated in Theorem 2. To make this requirement explicit, we compute the leading correction to lowest excited doubled-space eigenstate in the vectorized $K$-eigenbasis $\{|\epsilon_{p},\epsilon_{q}\rangle\}$,
@@
where $|\epsilon_{m}\rangle$ is the eigenstate of Lindblad operator $K$, $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pm}$ are the corresponding unperturbed doubled-space energy denominators. Requiring $\textrm{Tr}|\sigma_{1}|\leq\mathrm{poly}(L)$ then translates into a sparsity condition on $H$ in the $\{K_{j}\}$ eigenbasis:
$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|\epsilon_{s}\rangle,\qquad N(\alpha,H,|\epsilon_{s}\rangle)\leq\mathrm{poly}(L).$ (27)
Here $N(\alpha,H,|\epsilon_{s}\rangle)$ counts the number of indices $p$ for which $|H_{sp}|\geq e^{-\alpha L}$, with $|\epsilon_{s}\rangle$ a fixed eigenstate of $\{K_{j}\}$. The proof is given in the Appendix.
Weak-dissipation regime.— In the weak-dissipation regime we treat the dissipator as a perturbation to the unitary (Schrödinger) dynamics. We estimate (i) the Liouvillian gap from the leading dissipative correction to the unperturbed spectrum and (ii) the trace norm of the lowest excited state from the corresponding first-order correction to the eigenvectors. Combining these ingredients yields the mixing-time scaling.
Liouvillian gap. Consider an unperturbed doubled-space eigenstate
@@
where $\{|E_{m}\rangle\}$ are eigenstates of $H$. The leading (first-order) correction from the dissipative term is
$\langle E_{m},E_{n}^{(0)}|H_{d}|E_{m},E_{n}^{(0)}\rangle=-i\gamma\Bigl[\delta K_{m}^{2}+\delta K_{n}^{2}+\bigl(\overline{K_{m}}-\overline{K_{n}}\bigr)^{2}\Bigr],$ (28)
with $\overline{K_{m}}\equiv\langle E_{m}|K|E_{m}\rangle$ and
$\delta K_{m}^{2}\equiv\langle E_{m}|K^{2}|E_{m}\rangle-\overline{K_{m}}^{\,2}.$
@@
When $K,H]\neq 0$, typical fluctuations scale as $\delta K_{m}^{2}\sim L^{a}$ with $a\geq 0$. For gapped systems one expects $a=0$, giving a Liouvillian gap of order $\mathcal{O}(\gamma)$; for gapless systems ($a>0$), enhanced fluctuations can parametrically reduce the gap. [^5] If $[K,H]=0$, then $\delta K_{m}^{2}=0$ and decay rates are set by the smallest nonzero value of $(\overline{K_{m}}-\overline{K_{n}})^{2}$. For local $K$ in an ETH-satisfying system this scale is typically $\mathcal{O}(1)$, yielding a size-independent gap. Therefore, a sufficient condition for rapid mixing based on the Liouvillian gap is that $H$ is gapped, which ensures a constant gap.
@@
Trace norm of the lowest excited state. The first-order correction to the corresponding excited eigenvector takes the form
@@
$\begin{split}|E_{m},E_{n}^{(1)}\rangle=-i\gamma\sum_{p,q,s,t}\frac{\bigl(K_{ps}-K_{qt}\bigr)\bigl(K_{sm}-K_{tn}\bigr)}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle,\end{split}$ (29)
with $K_{ps}\equiv\langle E_{p}|K|E_{s}\rangle$. To ensure that the trace norm of the lowest excited state grows at most polynomially with system size, it is sufficient to impose a sparsity constraint on the matrix elements of $K$ in the $\{|E_{s}\rangle\}$ basis:
$|K_{sp}|\geq e^{-\alpha L}\quad\text{for fixed }|E_{s}\rangle.$
A simple limiting example is the nearest-neighbor quantum Ising model (with Hamiltonian containing only $\sigma_{j}^{z}\sigma_{j+1}^{z}$ terms), with Lindblad operators taken to be arbitrary single-site operators. In this case the Hamiltonian eigenbasis is a product basis, and each local $K$ connects only $\mathcal{O}(1)$ configurations. Consequently one may take $N(\alpha,K,|E_{s}\rangle)=\mathcal{O}(1)$ (e.g., $N=1$ in the strictly diagonal case), illustrating the required sparsity.
@@
Discussion— By using the double-space mapping, we directly connected the mixing time to the Liouvillian gap and the trace norm of the lowest excited states. This connection provides a clear requirement for achieving fast and rapid mixing. This framework also explains previous findings that mixing time in open systems depends on both the Liouvillian gap and the localization length of the eigenstates [29]. Our analysis shows that this is because localization properties directly affect the trace norm of the eigenstates, which in turn influences the mixing time.
We analyzed the mixing time in both the strong and weak dissipation regimes, where the Liouvillian gap and trace norm can be calculated straightforwardly using perturbation theory. This analysis not only simplifies the problem but also reveals general conditions for fast and rapid mixing. These findings are crucial for assessing the efficiency of quantum algorithms and state preparation using dissipative evolution. The general conditions we identified can be used to design specific forms of the Lindblad operator to more efficiently prepare desired quantum states.
@@
This work opens several interesting avenues for future research. First, it is important to compute the Liouvillian gap and the trace norm of the lowest excited states beyond the weak- and strong-dissipation limits, where our perturbative approach breaks down. Second, in connection with rapid mixing, it would be useful to identify conditions under which the trace norm of the lowest excited states is at most polynomial in system size, and to clarify how such conditions relate to features such as the locality of dissipation or the correlation length. A further question is whether changes in the scaling of the Liouvillian gap are tied to changes in the scaling of the lowest-mode trace norm. By tuning model parameters, one can compare the crossover where the inverse Liouvillian gap changes its system-size dependence (e.g., from polynomial to logarithmic) with the crossover where the trace norm changes (e.g., from exponential to polynomial). Determining whether these crossovers coincide—or how far apart they occur—could provide a new way to distinguish relaxation mechanisms. The boundary-driven transverse-field Ising chain, with dissipation only at the ends, provides a concrete setting in which both scalings vary with dissipation strength and can be directly compared. Finally, for dissipative state preparation, rapid mixing is essential to reach the target state on experimentally relevant timescales; the sparsity constraints derived here offer a practical guide for designing protocols that remain experimentally feasible.
@@
Acknowledgements— We thank Jerome Lloyd, Tian-Gang Zhou and Pengfei Zhang for helpful discussions.

### v2

@@ Abstract
A key concept in the study of open system dynamics is the mixing time, defined as the minimum time it takes for an arbitrary initial state to evolve to the system’s steady state [55, 82]. In dissipative state preparation, mixing time is a vital metric because it quantifies the practical time required to prepare a desired state, thus determining whether a protocol is experimentally feasible [19, 77]. Also, an estimation of the mixing time typically plays a crucial role in many of the quantum algorithmic applications to yield efficient quantum algorithms [29, 2]. Also, mixing time is related to entropy [32, 47, 14], entanglement dynamics, and the structure of the Lindblad spectrum [59, 60] and the dissipative phase transition, where at the transition point, the mixing time diverges [47, 82, 46, 43, 58, 34, 33].
Considerable progress has been made in bounding mixing times for open quantum dynamics. For Lindbladians satisfying quantum detailed balance, the dynamics can be mapped to a local frustration-free parent Hamiltonian on a doubled Hilbert space, whose spectral gap lower-bounds the Lindbladian gap and enables mixing-time estimates [16, 71, 36]. Complementarily, the modified logarithmic Sobolev inequality [76, 10] controls the exponential decay of relative entropy and can yield tighter, $\mathrm{polylog}(N)$ mixing-time bounds than gap-based estimates [37, 12, 5, 4, 24]. Other works bound mixing by treating the Hamiltonian and dissipative parts separately, but only in restricted settings [21]. Despite these advances, determining the mixing time and its system-size scaling for generic Lindbladians remains challenging and often becomes a bottleneck.
@@
That is, $\tau_{\mathrm{mix}}(\eta)$ is the minimal time such that, for arbitrary initial state $\rho_{0}$, the evolved state $e^{\mathcal{L}t}[\rho_{0}]$ is within trace distance $\eta$ of the steady state $\sigma$. Since density matrices are Hermitian, this distance can be written as
$D(\rho(t),\sigma)=\frac{1}{2}\,\mathrm{Tr}\,|\rho(t)-\sigma|.$
@@
Here $L$ and $R$ denote the left and right copies. Under this mapping, the Lindblad master equation becomes a Schrödinger-like equation in doubled space [79],
@@
$\begin{split}H_{s}=&H_{L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes H_{R}^{T},\\ H_{d}=&\gamma\sum_{\alpha}\Big[-2\,K_{\alpha,L}\otimes K_{\alpha,R}^{*}\\ &+\left(K_{\alpha}^{\dagger}K_{\alpha}\right)_{L}\otimes\mathcal{I}_{R}+\mathcal{I}_{L}\otimes\left(K_{\alpha}^{\dagger}K_{\alpha}\right)_{R}^{*}\Big].\end{split}$ (6)
@@
where $\epsilon_{j}=\alpha_{j}-i\beta_{j}$ are the eigenvalues of $H^{D}$ (the Lindblad spectrum). The real parts $\alpha_{j}$ generate oscillations, while $\beta_{j}\geq 0$ set decay rates  [^1] Complete positivity implies $\beta_{j}\geq 0$. Trace preservation fixes the steady mode: $c_{0}=1$, $\mathrm{Tr}\,\sigma_{0}=1$, and $\mathrm{Tr}\,\sigma_{j}=0$ for $j\neq 0$. Hermiticity preservation implies $\sigma_{j}=\sigma_{j}^{\dagger}$.. Consequently, the deviation from the steady state $\sigma_{0}$ is
$|\psi_{\rho}^{D}(t)\rangle-|\psi_{\sigma_{0}}\rangle=\sum_{j\neq 0}c_{j}\,e^{-i\alpha_{j}t-\beta_{j}t}\,|\psi_{\sigma_{j}}\rangle,$
@@
[…]
At time $t$, $|\psi_{\rho_{0}}(t)\rangle-|\psi_{\sigma_{0}}\rangle=c_{1}e^{-(\Delta-i\alpha_{1})t}\,|\psi_{\sigma_{1}}\rangle$. Consequently, the trace distance to the steady state is
$D(\rho(t),\sigma_{0})=\frac{|c_{1}|}{2}e^{-\Delta t}\operatorname{Tr}|\sigma_{1}|.$ (10)
@@
$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log\left(\operatorname{Tr}|\sigma_{1}|\right)-\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (11)
This reformulation of the mixing time indicates that the mixing time of an open system is set by the inverse Liouvillian gap, multiplied by a logarithmic factor controlled by the trace norm of the lowest excited state of the doubled-space Hamiltonian. One might worry that higher excited states, despite their larger decay rates, could have sufficiently smaller trace norms to yield a larger overall mixing-time bound. We explain in SM [1] why it is nevertheless sufficient to focus on the lowest excited state  [^2] We focus on the lowest excited state for two reasons. First, the fast mixing condition imposes no explicit constraints on trace norms across the spectrum, so the lowest excited state provides a natural and minimal diagnostic. Second, in both the weak- and strong-dissipation regimes, the rapid mixing requirement on the trace norm takes the same functional form for all excited states, and is therefore not specific to the lowest excited state. See SM [1] for details..
@@
$\tau_{\textit{mix}}(\eta)\leq\frac{1}{2}\Delta^{-1}\left[\log(N)-2\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (14)
@@
$\tau_{\text{mix}}(\eta)\leq\tfrac{1}{2}\Delta^{-1}\left[L\log d_{0}-2\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (15)
This mixing-time bound is general but need not be tight: its prefactor depends on the trace norm of the slowest mode, which can vary widely with the spatial structure of Liouvillian eigenmodes. If the slow mode is effectively supported on a subextensive region, $\operatorname{Tr}|\sigma_{1}|$ may scale subexponentially (or even $\mathrm{poly}(L)$), reducing the estimate from the generic scaling implied by $\operatorname{Tr}|\sigma_{1}|\sim e^{cL}$ to much smaller forms. Non-Hermitian skin effects that accumulate eigenstates near boundaries [30] offer a plausible route to such reduced support; we leave an explicit Lindbladian construction and a quantitative verification of the trace-norm scaling to future work. Overall, mixing is governed jointly by the Liouvillian gap and the spatial structure of the slow eigenmodes.
@@
We next apply these general criteria to the weak- and strong-dissipation regimes, where they simplify to explicit, checkable sufficient conditions expressed in terms of the Hamiltonian and Lindblad operators. The resulting rapid mixing conditions are summarized in Table 2.
@@
Strong-dissipation regime— In this regime, the Hamiltonian (Schrödinger) contribution can be treated perturbatively. We use degenerate perturbation theory to estimate the Liouvillian gap and the mixing time, and obtain the following fast-mixing corollary:
@@
$\langle\psi_{0k}^{L}|V^{D}|\psi_{1\alpha}^{R}\rangle\sim J,\qquad\langle\psi_{0k}^{L}|V^{D}|\psi_{0l}^{R}\rangle\sim J^{\prime},$
where $\{|\psi_{0k}^{L}\rangle\}$ spans the degenerate zeroth-order manifold (chosen to diagonalize $V^{D}$ within that subspace) and $\{|\psi_{1\alpha}\rangle\}$ denotes zeroth-order excited modes.
@@
Since the smallest diagonal shift $V^{D}_{0k,0k}$ is zero  [^4] In the vectorized representation, one has $V^{D}=H\otimes\mathbb{I}-\mathbb{I}\otimes H^{T}$, which acts as the commutator superoperator, $V^{D}\,|\rho\rangle\rangle=|[H,\rho]\rangle\rangle$. Therefore, any operator $\rho$ that commutes with $H$ is a zero mode of $V^{D}$. In particular, the identity operator (and more generally any function of $H$, or any $|m\rangle\langle m|$ in the energy basis) satisfies $H,\rho]=0$, implying an exact eigenvalue $\lambda=0$ of $V^{D}$. Since our $\{|\psi_{0k}^{R}\rangle,\langle\psi_{0k}^{L}|\}$ are chosen to diagonalize $V^{D}$ within the zeroth-order degenerate manifold, the diagonal matrix elements $\langle\psi_{0k}^{L}|V^{D}|\psi_{0k}^{R}\rangle$ coincide with these eigenvalues; hence the smallest value by magnitude is necessarily $\min_{k}|\lambda_{k}|=0$., thus the Liouvillian gap is set by the dissipative contribution from the second-order term,
$\Delta\sim\sum_{q}\frac{|V^{D}_{0k,1q}|^{2}}{E^{(0)}_{1q}-E^{(0)}_{0k}}\sim c\,\frac{J^{2}}{\gamma_{\mathrm{loc}}},$ (20)
with an $\mathcal{O}(1)$ constant $c$ (for our boundary coupling structure, c=8c=8) [^5] The factor $8$ counts: (i) action on the left or right copy, (ii) action at the left or right boundary, and (iii) raising or lowering the boundary state, giving $2^{3}=8$ possibilities.. Crucially, $\Delta$ does not shrink with $L$: only Hamiltonian terms involving boundary sites connect the steady manifold to decaying modes, so the number of contributing excited states is set by the boundary size, i.e., $\mathcal{O}(1)$ in 1D.
@@
$\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}\log\!\bigl(\mathrm{Tr}|\sigma_{1}|\bigr).$
@@
and this bound is saturated here because the $V^{D}$-diagonal basis typically forms superpositions across the full manifold. Therefore, $\tau_{\mathrm{mix}}(\eta)\sim\frac{\gamma_{\mathrm{loc}}}{c\,J^{2}}(\log d_{0})\,L$, which is $\mathrm{poly}(L)$ and hence implies fast mixing. In the SM [1], we analyze the strong bulk-dissipation regime and show that this perturbative mechanism does not apply.
Physically, boundary-driven fast mixing is a quantum-Zeno-type stabilization: strong edge dissipation sets an $\mathcal{O}(1)$ scale, while boundary Hamiltonian couplings generate a size-independent gap $\Delta\propto J^{2}/\gamma_{\mathrm{loc}}$. Thus, boundary dissipation can guarantee scalable mixing even for complex bulk dynamics, whereas bulk dissipation typically requires different control assumptions (see SM [1]).
Finally, in the strong-dissipation regime, rapid mixing also requires control of the trace norm $\mathrm{Tr}\,|\sigma_{1}|$, as stated in Theorem 2. To make this requirement explicit, we assume that $K$ is diagonalizable and compute the leading correction to the lowest excited doubled-space eigenstate in the vectorized $K$-eigenbasis $\{|\epsilon_{p},\epsilon_{q}\rangle\}$,
@@
where $\{|\epsilon_{m}\rangle\}$ is the eigenstate of Lindblad operator $K$, $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pm}$ are the corresponding unperturbed doubled-space energy denominators. Requiring $\textrm{Tr}|\sigma_{1}|\leq\mathrm{poly}(L)$ then translates into a sparsity condition on $H$ in the $\{K_{j}\}$ eigenbasis:
$\exists\,\alpha\geq 1\ \text{s.t.}\ \forall\,|\epsilon_{s}\rangle,\qquad N(\alpha,H,|\epsilon_{s}\rangle)\leq\mathrm{poly}(L).$
Here $N(\alpha,H,|\epsilon_{s}\rangle)$ counts the number of indices $p$ for which $|H_{sp}|\geq e^{-\alpha L}$ (with $s$ fixed). The proof is given in the SM [1].
Weak-dissipation regime—In the weak-dissipation regime, we treat the dissipator as a perturbation to the unitary (Schrödinger) dynamics. We estimate (i) the Liouvillian gap from the leading dissipative correction to the unperturbed spectrum and (ii) the trace norm of the lowest excited state from the corresponding first-order correction to the eigenvectors. Combining these ingredients yields the mixing-time scaling.
Liouvillian gap. Consider an unperturbed doubled-space eigenstate with eigenvalue $\epsilon_{mn}^{(0)}$:
@@
where $\{|E_{m}\rangle\}$ are the eigenstates of $H$. The first-order correction from the dissipative term is
$\langle E_{m},E_{n}^{(0)}|H_{d}|E_{m},E_{n}^{(0)}\rangle=-i\gamma\Bigl[\delta K_{m}^{2}+\delta K_{n}^{2}+\bigl(\overline{K_{m}}-\overline{K_{n}}\bigr)^{2}\Bigr]$
@@
with $\overline{K_{m}}\equiv\langle E_{m}|K|E_{m}\rangle$ and $\delta K_{m}^{2}\equiv\langle E_{m}|K^{2}|E_{m}\rangle-\overline{K_{m}}^{\,2}$. When $K,H]\neq 0$, typical fluctuations scale as $\delta K_{m}^{2}\sim L^{a}$ with $a\geq 0$. For gapped systems one expects $a=0$, giving a Liouvillian gap of order $\mathcal{O}(\gamma)$; for gapless systems ($a>0$), enhanced fluctuations can parametrically reduce the gap  [^6] If $[K,H]=0$, then $\delta K_{m}^{2}=0$ and decay rates are set by the smallest nonzero value of $(\overline{K_{m}}-\overline{K_{n}})^{2}$. For local $K$ in an ETH-satisfying system, this scale is typically $\mathcal{O}(1)$, yielding a size-independent gap.. Therefore, a sufficient condition for rapid mixing based on the Liouvillian gap is that $H$ is gapped, which ensures a constant gap.
@@
Trace norm of the lowest excited state. The first-order correction to the corresponding excited eigenvector is
@@
$\begin{split}|E_{m},E_{n}^{(1)}\rangle=-i\gamma\sum_{pqst}\frac{\bigl(K_{ps}-K_{qt}\bigr)\bigl(K_{sm}-K_{tn}\bigr)}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle,\end{split}$ (22)
[…]
@@
Discussion— In this letter, by using the double-space mapping, we directly connected the mixing time to the Liouvillian gap and the trace norm of the lowest excited states. This connection provides a clear requirement for achieving fast and rapid mixing. This framework also explains previous findings that mixing time in open systems depends on both the Liouvillian gap and the localization length of the eigenstates [30]. Our analysis shows that this is because localization properties directly affect the trace norm of the eigenstates, which in turn influences the mixing time. We analyze the mixing time in both the strong and weak dissipation regimes, where the Liouvillian gap and trace norm can be calculated straightforwardly using perturbation theory. This analysis not only simplifies the problem but also reveals general conditions for fast and rapid mixing.
@@
This work suggests several directions for future research. First, motivated by rapid mixing, it would be valuable to identify general conditions under which the trace norm of the slowest mode is bounded by $\mathrm{poly}(L)$, and to understand how these conditions relate to physical features such as locality of dissipation and correlation length. A closely related question is whether changes in the scaling of the Liouvillian gap are accompanied by changes in the scaling of the slowest-mode trace norm. Concretely, one can tune model parameters to compare (i) the crossover where the inverse gap changes its system-size dependence (e.g., from polynomial to logarithmic) with (ii) the crossover where the trace norm changes (e.g., from exponential to polynomial). Whether these crossovers coincide could provide a diagnostic of distinct relaxation mechanisms. The boundary-driven transverse-field Ising chain, with dissipation only at the ends, offers a concrete setting where both scalings vary with dissipation strength and can be directly compared. Finally, rapid mixing is crucial for dissipative state preparation on experimentally relevant timescales. Our sparsity constraints provide practical guidance for designing experimentally feasible protocols and for assessing the efficiency of dissipative algorithms. More broadly, the general conditions identified here can be leveraged to engineer Lindblad operators that prepare desired target states more efficiently.
@@
Acknowledgements— We thank Jerome Lloyd, Tian-Gang Zhou, Pengfei Zhang, Yu-Min Hu, Shuo Liu, and Yuxuan Zhang for helpful discussions. We are also grateful to Lorenza Viola and Yikang Zhang for their careful reading of the manuscript and for providing valuable feedback and suggestions.